(R=0.082 L atm mol-1 K-1)
Calculate the absolute temperature of 21 g Nitrogen gas at which gas occupies 164 mL at 6 atm.
1
Expert's answer
2020-06-23T02:17:05-0400

"This\space problem\space is\space solved\space using\space the\space Ideal\space Gas\space Law\space formula.\newline
\space \space\space PV=nRT\newline n\space is\space number\space of\space moles\space which\space is=\dfrac{21g}{28.02}"


"=0.75mol"

"Convert\space volume(164ml)\space into\space litres(L)\newline\space \space \space \dfrac{164}{1000}=0.164L\newline Rearraging\space the\space formula\space we \space get,\newline\space \space \space T=\dfrac{PV}{nR} \newline \newline \space \space \space T=\dfrac{6*0.164}{0.75*0.082}=16\newline The\space absolute\space temperature\space is\space16K"